OpenStudy (anonymous):

I need help...I don't understand or know what to do or where to start. F(x)=-2(x^2-4)/3(x^2+4x+4)

jhonyy9 (jhonyy9):

2(x-2)(x+2) ----------- = 2(x-2)/(3(x+2)) 3(x+2)^2 - hope so much that is understandably sure

jhonyy9 (jhonyy9):

good luck bye

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

so it is just pretty much like an elimination thing

jhonyy9 (jhonyy9):

yes right

OpenStudy (anonymous):

so do i try to solve the final equation or no

jhonyy9 (jhonyy9):

x^2 -4 =(x-2)(x+2) --- us formula a^2 -b^2 =(a-b)(a+b) x^2 +4x +4 =(x+2)^2

jhonyy9 (jhonyy9):

you need to solve it again ?

OpenStudy (anonymous):

so on the 1st 1 would it be 0--x cancels and 2 cancels

jhonyy9 (jhonyy9):

so sorry i have missed yes there is -2(x-2)/(3(x+2))

OpenStudy (anonymous):

im not understanding what you were saying there in your last response

jhonyy9 (jhonyy9):

before i have wrote there in numerator 2(x-2/ i now i have said that there is -2

jhonyy9 (jhonyy9):

so you need to solve it ?

jhonyy9 (jhonyy9):

than you need to suppose that x not can being -2 because than the denominator will be equal zero and than the fraction is undefined

jhonyy9 (jhonyy9):

but x=2 is right

jhonyy9 (jhonyy9):

like solution of it

OpenStudy (anonymous):

yes i do understand now thank you

jhonyy9 (jhonyy9):

yw my pleasure good luck bye